OpenStudy (anonymous):

write the slope-intercept form of the equation of the line that passes through the point (-5,4) and has a slope of -1

OpenStudy (anonymous):

Slope-intercept form of the line is y = mx + b where m is the slope and b is the y-intercept Using the Point-Slope form we first get the equation and then convert it to slope-intercept form point (-5,4) and has a slope of -1 so m = -1, x coordinate of the given point = -5 and y-coordinate = 4 Point-Slope form : y - y1 = m(x - x1) substituting given values we get y - 4 = -1[x - (-5)] y - 4 = -1[x + 5] y - 4 = -x -5 y + x = -5 + 4 y - x = -1 putting it in the Slope-Intercept form, we get y = x - 1